You and your friend both having equal masses of 50kg are seperated by a mass of 1m.calculate the force of gravitation between you and your friend
1
Expert's answer
2019-09-23T09:12:01-0400
"F=\frac{Gm^2}{r^2}"

"F=(6.67\cdot 10^{-11})\frac{50^2}{1^2}=1.67\cdot 10^{-7}N"